Jacque is twice as old as her sister Chezi. When Jacque is twice as old as she is now, she will be six years older than Chezi. How old is Jacque now?
Answer by macston(5194)

If Jacque will be 6 years older than Chezi at any time, it means she is always 6 years older than Chezi.
.
J=Jacque's age; C=Chezi's age
.
J=C+6
J-6=C
.
J=2C
J=2(J-6)
J=2J-12
12=J ANSWER: Jacque is 12 years old.